1) Daniel is planning to earn money over the summer mowing lawns in the neighborhood to buy a scooter. The scooter he wants costs $2,200. He will be charging $20/lawn and estimates his costs to be $2/lawn. How many lawns must he mow in order to pay for the scooter entirely through his earnings?

*show your work*
Answer #1

Cost of the scooter =2200

Price charged per lawn =20

Cost per lawn =2

Profit per lawn=Price charged -cost =20-2=$ 18

Number of lawns Daniel must mow =Cost of the scooter/ Profit per lawn

=2200/18

=122.22

=123 lawns (rounding to next whole number as we assume that portion of lawn cannot be mowed)